OpenStudy (anonymous):

Simplify: the square root of 6 (7 the square root of 3 + 6)

OpenStudy (anonymous):

the square root of 6 is 3 and the square root of 6=3 is 3.

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

cant understand the question :s

OpenStudy (anonymous):

Are there any brackets in this? For example is it \(\sqrt{6}(7\sqrt{3}+6)\) Or is it\(\sqrt{6}(7\sqrt{3+6})\)

OpenStudy (anonymous):

You are correct hea

OpenStudy (anonymous):

The first one is right